# Random Walk in Random Permutation Set Theory

### v1

@@ Abstract
Shannon entropy, Deng entropy, Evidence theory, Approximation, Random permutation set, Maximum entropy
@@ I Introduction
Random walk models have been used to simulate various natural processes. These models are particularly useful for understanding molecular-level dynamics (Kessing et al., 2022; Ansari-Rad, Abdi, and Arzi, 2012; Thompson, Kienle, and Schwartz, 2022; Ding, Zhang, and Liu, 2023), graphical analysis (Lin, Julaiti, and Zhang, 2012; Zhu et al., 2023) and so on (Yang et al., 2022; Cherayil, 2024). Correlated walks are specialized category of random walks. In correlated walks, the moving particles have a memory of their previous steps. This memory affects the direction of the next step, making the order of the steps important (Tojo and Argyrakis, ).
The Dempster-Shafer evidence theory (DSET), also known as evidence theory, is a framework used for reasoning under uncertainty (AP, 1967; Shafer, 1976). In contrast to probability theory, DSET utilizes mass functions to assign beliefs to subsets of a power set, rather than to individual outcomes in the sample space, allowing for a more flexible approach to defining belief assignments. Thus, DSET has been applied to various fields (Liu, Zhao, and Xiao, 2023; Gao et al., 2023; Liu et al., 2024; Yang and Xiao, 2024) However, DSET struggles to handle ordered information in certain real-world problems. To address this limitation, the random permutation set theory(RPST) is introduced (Deng, 2022). By introducing the concept of permutation events, RPST effectively considers the order of elements and expands the power set and mass function into the permutation events space (PES) and permutation mass function (PMF). Similar to Shannon entropy (Shannon, 1948) in probability and Deng entropy (Deng, 2016) in DSET, the RPS entropy is proposed by Chen and Deng (2023) to quantify the uncertainty in RPST.
One area of focus in RPST research pertains to its physical implications. For instance, Zhan, Li, and Deng (2024) delved into expanding order information to encompass more complex relationships. Zhou et al. (2022) used cooperative game to interpret RPST. Deng and Deng (2022) discovered the PMF condition for maximum entropy in RPST. And Zhao, Li, and Deng (2023) further demonstrated that the information dimension associated with this PMF condition is $2$, akin to the fractal dimension of Brownian motion. Moreover, the mean square distance, denoted as $\bar{r^{2}}$, in a Brownian motion is proportional to the time elapsed. In our previous research, it is showed that the limit form of maximum RPS entropy is $e\cdot(n!)^{2}$, exhibiting similarities to $\bar{r^{2}}$ (Zhou et al., 2024). These collective discoveries hint at a potential connection between RPST and Brownian motion, or random walk in mathematics.
@@ II.2.2 RPS entropy
RPS entropy is fully compatible with Deng entropy (Deng, 2016) as used in evidence theory, and Shannon entropy (Shannon, 1948) in probability theory. Such uncertainty measure has been provided insights for other uncertainty measures like distance (Chen, Deng, and Cheong, 2023a), divergence (Chen, Deng, and Cheong, 2023b), information fractal dimension (Zhao, Li, and Deng, 2023) and so on.

### v2

@@ Abstract
Random Permutation Set, Random walk, Probability theory, Shannon entropy, Brownian motion, Wiener process
@@ I Introduction
Random walk models have been used to simulate various natural processes. These models are particularly useful for understanding molecular-level dynamics (Kessing et al., 2022; Ansari-Rad, Abdi, and Arzi, 2012; Thompson, Kienle, and Schwartz, 2022), complex networks (Zhang, Li, and Deng, 2018; Wang et al., 2021), and so on. Correlated walks are specialized category of random walks. In correlated walks, the moving particles have a memory of their previous steps. This memory affects the direction of the next step, making the order of the steps important (Tojo and Argyrakis, 1996).
The Dempster-Shafer evidence theory (DSET), also known as evidence theory, is a framework used for reasoning under uncertainty (AP, 1967; Shafer, 1976). In contrast to probability theory, DSET utilizes mass functions to assign beliefs to subsets of a power set, rather than to individual outcomes in the sample space, allowing for a more flexible approach to defining belief assignments. Thus, DSET has been applied to various fields (Liu, Zhao, and Xiao, 2023; Gao et al., 2023; Liu et al., 2024; Yang and Xiao, 2024; Contreras-Reyes and Kharazmi, 2023; Cui et al., 2022) However, DSET struggles to handle ordered information in certain real-world problems. To address this limitation, the random permutation set theory(RPST) is introduced (Deng, 2022). By introducing the concept of permutation events, RPST effectively considers the order of elements and expands the power set and mass function into the permutation events space (PES) and permutation mass function (PMF). Similar to Shannon entropy (Shannon, 1948) in probability and Deng entropy (Deng, 2016) in DSET, the RPS entropy is proposed by Chen and Deng (2023) to quantify the uncertainty in RPST.
One research area of focus in DSET and RPST is its physical implications. For instance, in DSET Li and Xiao (2023) derived normal distribution from maximum Deng entropy Kang and Deng (2019), and Zhao, Li, and Deng (2024) found intriguing linearity in Deng entropy. In RPST, Zhan, Li, and Deng (2024) expanded order information to encompass more complex relationships. Zhou et al. (2022) used cooperative game to interpret RPST. Deng and Deng (2022) discovered the PMF condition for maximum entropy in RPST. And Zhao, Li, and Deng (2023) further demonstrated that the information dimension associated with this PMF condition is $2$, similar to the fractal dimension of Brownian motion. Moreover, the mean square distance, denoted as $\bar{r^{2}}$, in a Brownian motion is proportional to the time elapsed. In our previous research, it is showed that the limit form of maximum RPS entropy is $e\cdot(n!)^{2}$, exhibiting similarities to $\bar{r^{2}}$ (Zhou et al., 2024). These collective discoveries hint at a potential connection between RPST and Brownian motion, or random walk in mathematics.
@@ II.2.2 RPS entropy
RPS entropy is fully compatible with Deng entropy (Deng, 2016) as used in evidence theory, and Shannon entropy (Shannon, 1948) in probability theory. Such uncertainty measures have been provided insights for other uncertainty measures like distance (Chen, Deng, and Cheong, 2023a), divergence (Chen, Deng, and Cheong, 2023b; Zeng and Xiao, 2023), information measures (Zhao, Li, and Deng, 2023; Kharazmi and Contreras-Reyes, 2023; Ortiz-Vilchis, Lei, and Ramirez-Arellano, 2024) and so on.